# Applications of Spiking Neural Networks in Visual Place Recognition

### v2

@@ Abstract
In robotics, Spiking Neural Networks (SNNs) are increasingly recognized for their largely-unrealized potential energy efficiency and low latency particularly when implemented on neuromorphic hardware. Our paper highlights three advancements for SNNs in Visual Place Recognition (VPR). Firstly, we propose Modular SNNs, where each SNN represents a set of non-overlapping geographically distinct places, enabling scalable networks for large environments. Secondly, we present Ensembles of Modular SNNs, where multiple networks represent the same place, significantly enhancing accuracy compared to single-network models. Each of our Modular SNN modules is compact, comprising only 1500 neurons and 474k synapses, making them ideally suited for ensembling due to their small size. Lastly, we investigate the role of sequence matching in SNN-based VPR, a technique where consecutive images are used to refine place recognition. We analyze the responsiveness of SNNs to ensembling and sequence matching compared to other VPR techniques. Our contributions highlight the viability of SNNs for VPR, offering scalable and robust solutions, and paving the way for their application in various energy-sensitive robotic tasks.
@@ I Introduction
Spiking Neural Networks (SNNs) represent a cutting-edge paradigm in neuromorphic computing, mirroring the intricate workings of biological neural systems [1, 2, 3, 4]. In these networks, every neuron possesses its distinct activation state. Unlike conventional neural networks, where neuron activations are typically continuous values, neurons in SNNs convey information through intermittent spikes, which are initiated when the neuron’s activation surpasses a particular threshold [1, 5, 6]. These spiking networks exhibit promising attributes when deployed on neuromorphic hardware, offering notable energy efficiency and low-latency data processing [7, 8, 2, 9, 10]. Despite these potential advantages, SNNs have seen minimal adaptation in robotics, due to limitations such as difficulty in supervised training of SNNs due to the non-differentiable activation function of spiking neurons and a lack of tools and resources [2, 11, 6, 12, 4].
@@
Visual place recognition is a critical component in robot localization tasks such as loop closure detection in SLAM, and global re-localization of mobile robots [19, 13, 18, 16]. It is also relevant to image retrieval and landmark recognition tasks [15, 13]. Within robot navigation, VPR can minimize localization errors by recognizing previously visited places and updating the map of the environment despite appearance changes [15, 13, 20], which enables mobile robots to operate over extended periods.
To operate in real-time on resource-constrained robots, such as space exploration and disaster recovery where long mission times are desirable, conventional SOTA VPR methods might not be applicable as they often have high computational demands [21], motivating the use of SNNs within VPR. We are further inspired by the remarkable ability of animals with relatively small brains, such as rodents, to effectively perform navigation in complex environments. Although SNNs are not as mature as their deep learning counterparts, the benefits of SNNs include low power usage and low latency, particularly when deployed on neuromorphic hardware [8, 2, 3]. VPR is a particularly intriguing task for SNNs as VPR is amenable to enhancement strategies such as ensembling and using temporal data via sequences of images.
@@
Ensembling [^1] We note ensembling [29] and model fusion [30] have been used interchangeably in literature to denote the practice of combining multiple models or feature representations for improved performance. Throughout this work, we refer to this technique as ensembling. is the strategy of combining multiple models to boost accuracy, reduce overfitting, and yield more robust models [1, 29, 31, 32, 33, 34]. While there is an inherent trade-off between the number of ensemble members and the energy efficiency of a system [29, 31], the benefits of ensembling, particularly its enhanced accuracy and reliability, are invaluable to navigation-related robotics applications. Ensembling has been previously utilized in the context of monocular SLAM [35] for improved data association, terrain segmentation for learning data collected at different times [36], and place recognition with image-based [37, 38] and event-based [39] input, where multiple VPR techniques fuse to enhance place recognition robustness. Similarly, our work also utilizes ensembling to improve upon the robustness and generalization ability of our system but differs with these works in terms of how we define our ensemble members. Our ensemble members all are tasked to do the same task, have the same architecture and differ in random shuffling of input images and initial learned weights. A similar approach of employing ensembles was previously demonstrated in uncertainty estimation to provide sufficient diversity among the members, and improve the overall performance of the system [40].
In the context of VPR, a sequence matching technique uses consecutive reference frames instead of single frames, to match a query image of a place to its corresponding reference image [41, 42, 43]. By analyzing a series of images, this technique enables enhanced resilience against temporary environmental disruptions and improves localization accuracy [41, 42, 43]. Sequence matching techniques have been widely employed in place recognition [41, 44, 43, 45, 42, 46, 47], where often a decoupled approach consisting of an initial image-based retrieval method and subsequent sequence score aggregation [41, 44, 47] is employed. Similar to these techniques, our paper utilizes sequence matching as an additional step after image retrieval. We further demonstrate the effectiveness of sequence matching for SNN-based approaches, and provide an indicator for responsiveness of conventional and SNN-based VPR methods to sequence matching.
@@ II Related works
In this section, we offer an overview of neuromorphic computing and spiking neural networks (Section II-A) and explore the applications of spiking neural networks (SNNs) in robot localization (Section II-B). We then delve into the visual place recognition task (Section II-C), non-spiking biologically inspired VPR approaches (Section II-D) and provide insights into ensembling (Section II-E) and sequence matching techniques (Section II-F).
@@ II-B Spiking Neural Networks for Robot Localization
The capabilities of SNNs have been demonstrated in a wide range of computer vision and robotics tasks. These include pattern recognition [73], control [74, 75, 76, 77, 78], manipulation [79, 80, 81], object tracking [82, 83], and scene understanding [84, 85]. Many works have employed SNNs to address tasks related to robot localization, which is the problem under investigation in this work. These works include RatSLAM [86] which computationally models place, grid and border cells of the rat hippocampus[87], a navigation controller for mapping unknown environments [88], a pose estimation and map formation method [89], a light-weight system for uni-dimensional SLAM [90], and a SLAM model that utilizes representations of continuous spatial maps to produce compressed structures from multiple domains [91]. In previous work [59], we presented a SNN specifically for VPR. This network had a limited capacity, recognizing up to just 100 places.
While some of these systems have also been deployed on neuromorphic hardware [87, 89, 90, 92, 93], so far, the performance of SNN-based methods for robot localization have only been demonstrated in simulated environments [88, 89, 91], constrained indoor environments [87, 90, 92, 93, 94], or small-scale outdoor environments [59].
@@ II-C Visual Place Recognition
The visual place recognition (VPR) task is to determine whether a place has been previously visited, even when faced with appearance changes and perceptual aliasing [15, 17, 18, 14, 13, 20]. VPR is most-commonly framed as an image retrieval problem, where feature representations of a given query image is compared to the feature representations of all previously visited places, i.e. the reference database. The predicted place of the query image is the true position of the reference place that is most similar to the query in feature space. The VPR problem can also be posed as a template matching problem, similar to an image classification problem, where associated templates of all reference images are extracted to represent each place either via a single image [100, 49, 101, 48] or a sequence of images [41, 42, 102].
@@ II-E Ensembles of Neural Networks
One well-known approach to improve the predictive performance of neural networks is to use an ensemble of models [1, 29, 31, 32, 33, 34]. Ensembles have been shown to generalize well and prevent issues such as over-fitting and instability, which makes these approaches suitable to a wide range of applications in different domains [1, 29, 31, 32, 33, 34]. Challenges in deploying ensembles include requiring sufficient diversity in the output of the individual ensemble members, the trade-off between the computational complexity and the number of ensemble members, and the predictive performance and latency of the ensemble [1, 29, 31, 32, 33, 34]. Although ensembles are typically limited in scalability, our work anticipates leveraging neuromorphic computing, renowned for its highly efficient parallelization capabilities [8], which alleviates scalability concerns in our ensembling approach due to the small size of each individual network.
Ensemble techniques have been used for a wide variety of robotic applications including image segmentation [114], gaze estimation [115], and uncertainty estimation [40]. In the field of SNNs, a variety of ensemble methods have been applied to diverse pattern recognition tasks [116, 117, 118, 119, 120, 121]. These methods include unsupervised ensembles of spiking expectation maximization networks [116], and ensembles of evolutionary SNN algorithms [119] that have been effective in digit recognition. Additionally, there are heterogeneous ensembles for few-shot online learning [117], and SNN ensembles that utilize a convolutional structure with unsupervised STDP learning [118], and ensembles of Liquid State Machines [121] which have been applied to image [117, 118, 121] and audio recognition [121] tasks. Furthermore, reservoir computing ensembles have been explored for multi-object behavior recognition [120].
@@ II-F Sequence Matching Techniques for Neural Networks
One common approach to improve the robustness of a VPR method, especially against sudden high appearance changes and perceptual aliasing, involves utilizing the temporal information inherent in the database and query sets used in the mobile robot localization task [17, 15, 14, 13, 20]. One such set of algorithms that use the temporal information are sequence matching techniques, which can be classified into similarity-based methods, feature-based methods, and approaches that learn sequential information [13].
@@ III-A Preliminaries
Figure 2: SNN Module Architecture: Our Modular SNN is comprised by SNN modules that all have the same three-layer SNN architecture illustrated in this figure. Each module converts an input image to spike trains via rate coding, where the firing rate of input neurons is based on pixel intensities. The total number of input neurons is equal to the number of pixels in the input image, denoted as $K_{P}=W\times H$. These input neurons (blue dots) are fully connected to a layer of excitatory neurons (blue arrows connecting to green dots). Each excitatory neuron is connected to a single inhibitory neuron (single green arrow connecting to a red dot), which in turn connects to and inhibits all other excitatory neurons except its paired excitatory neuron (red arrows connecting back to green dots). The synaptic weights from excitatory to inhibitory neurons, $W_{EI}$, and from inhibitory back to excitatory neurons, $W_{IE}$, are fixed constants. The synaptic weights from input neurons to excitatory neurons, $W_{PE}$, is learned via the unsupervised Spike-Time Dependent Plasticity (STDP) mechanism that enables excitatory neurons to respond to different places. The number of output spikes from these excitatory neurons is used for place predictions.
@@ III-C Hyperactive Neuron Detection
The basic fusion approach, that considers all spiking neurons of all modules as presented in Equation 7, is problematic. As the modules are only ever exposed to their local subset of the training data, there is a lack of global regularization to unseen training data outside of their local subset. In the case of spiking networks, this phenomenon leads to “hyperactive” neurons that are spuriously activated when stimulated with images from outside their training data. We decided to detect and remove these hyperactive neurons. This global regularization enhances place recognition capability within the modularity technique, allowing mapping to a larger number of places.
@@ III-D Ensemble of Modular SNNs
$D_{\text{single}}(u,v)=\max(S_{\text{single}})-S_{\text{single}}(u,v)$ (12)
@@ III-E Sequence Matching
where $I_{L_{seq}}$ is an identity matrix acting as a square filter kernel with dimensions of $L_{seq}\times L_{seq}$ pixels.
@@ IV-A Implementation Details
We implement our SNNs using the Brian2 simulator. We publicly release our code at: https://github.com/QVPR/VPRSNN. Each SNN module contains $K_{E}=400$ excitatory neurons, $K_{I}=400$ inhibitory neurons, and $K_{P}=W\times H$ input neurons, where the input image height and width is $28$ pixels in each dimension. We select the network’s hyperparameters as follows: For the number of training epochs, we use a fixed value of $30$ epochs for all datasets. For the threshold value $\theta$ (Eq. (8)) to detect the hyperactive neurons for each individual Modular SNN evaluated on a dataset, we draw the threshold value from a uniformly random distribution with $U(40,100)$ for small-scale datasets ($<1000$ places) and $U(600,800)$ for large-scale datasets ($>=1000$ places). The threshold value of each Modular SNN within an Ensemble of Modular SNNs evaluated on a dataset, is chosen separately based on this uniformly random selection process.
@@ IV-B Datasets
The Oxford RobotCar dataset [55]: has over 100 traverses captured in Oxford city and it is recorded at varying conditions including different time of the day, and different seasons. As done in prior works [131, 59], we selected the front left stereo frames from the Rain (2015-10-29-12-18-17) traverse as the reference dataset and from the Dusk (2014-11-21-16-07-03) traverse as the query dataset. For each traverse, we considered every image as a distinct place, gaining 3800 places.
The SFU-Mountain dataset [56]: has more than 8 hours of trail driving in Burnaby Mountain, British Columbia Canada, using the Clearpath Husky robot covering sunny, rainy and snowy conditions [56]. Following prior works [132, 133], we used the front right stereo frames from the dry traverse for reference and from the dusk traverse for query. We considered each image in the traverse as a different place, and used the author’s place sampling configurations [56], utilizing the entire dataset where each traverse contains 375 places.
The Synthia Night-to-Fall dataset [57]: is a synthetic dataset that was initially designed for semantic scene understanding in a city-like driving scenario. In our approach, following the methodology described in [134], we utilized the SEQS-04 foggy (reference) and nighttime (query) traverses. Segments in which the vehicle remained stationary were excluded, and we sampled approximately every 1 frame per second resulting in 250 places for each traverse.
The St Lucia dataset [58]: comprises several traverses along a route within the St Lucia suburb of Brisbane. For our experiments, we employed a traverse conducted during the early morning (190809-0845) as the reference traverse and another in the late afternoon (180809-1545) as the query traverse. We omitted the segments where the vehicle was at rest and sampled places approximately every 15 meters. We used only the unique places from the reference traverse, obtaining 500 places. For the query traverse, we included sections where places were visited multiple times within the same traverse, obtaining 1037 places.

### v3

@@ Abstract
In robotics, Spiking Neural Networks (SNNs) are increasingly recognized for their largely-unrealized potential energy efficiency and low latency particularly when implemented on neuromorphic hardware. Our paper highlights three advancements for SNNs in Visual Place Recognition (VPR). Firstly, we propose Modular SNNs, where each SNN represents a set of non-overlapping geographically distinct places, enabling scalable networks for large environments. Secondly, we present Ensembles of Modular SNNs, where multiple networks represent the same place, significantly enhancing accuracy compared to single-network models. Each of our Modular SNN modules is compact, comprising only 1500 neurons and 474k synapses, making them ideally suited for ensembling due to their small size. Lastly, we investigate the role of sequence matching in SNN-based VPR, a technique where consecutive images are used to refine place recognition. We demonstrate competitive performance of our method on a range of datasets, including higher responsiveness to ensembling compared to conventional VPR techniques and higher R@1 improvements with sequence matching than VPR techniques with comparable baseline performance. Our contributions highlight the viability of SNNs for VPR, offering scalable and robust solutions, and paving the way for their application in various energy-sensitive robotic tasks.
@@ I Introduction
Spiking Neural Networks (SNNs) represent a cutting-edge paradigm in neuromorphic computing, mirroring the intricate workings of biological neural systems [1, 2, 3, 4]. In these networks, every neuron possesses its distinct activation state. Unlike conventional neural networks, where neuron activations are typically continuous values, neurons in SNNs convey information through intermittent spikes, which are initiated when the neuron’s activation surpasses a particular threshold [1, 5, 6]. These spiking networks exhibit promising attributes when deployed on neuromorphic hardware, offering notable energy efficiency and low-latency data processing [7, 8, 2, 9, 10]. Despite these potential advantages, SNNs have seen minimal adaptation in robotics, due to limitations such as the difficulty of supervised training caused by the non-differentiable activation function of spiking neurons and a lack of tools and resources [2, 11, 6, 12, 4].
@@
Visual place recognition is a critical component in robot localization tasks such as loop closure detection in Simultaneous Localization and Mapping (SLAM), and global re-localization of mobile robots [19, 13, 18, 16]. It is also relevant to image retrieval and landmark recognition tasks [15, 13]. Within robot navigation, VPR can minimize localization errors by recognizing previously visited places and updating the map of the environment despite appearance changes [15, 13, 20], which enables mobile robots to operate over extended periods.
To operate in real-time on resource-constrained robots, such as space exploration and disaster recovery where long mission times are desirable, conventional state-of-the-art VPR methods might not be applicable as they often have high computational demands [21], motivating the use of SNNs within VPR. We are further inspired by the remarkable ability of animals with relatively small brains, such as rodents, to effectively perform navigation in complex environments. Although SNNs are not as mature as their deep learning counterparts, the benefits of SNNs include low power usage and low latency, particularly when deployed on neuromorphic hardware [8, 2, 3]. VPR is a particularly intriguing task for SNNs as VPR is amenable to enhancement strategies such as ensembling and using temporal data via sequences of images.
@@
Ensembling [^1] We note ensembling [29] and model fusion [30] have been used interchangeably in literature to denote the practice of combining multiple models or feature representations for improved performance. Throughout this work, we refer to this technique as ensembling. is the strategy of combining multiple models to boost accuracy, reduce overfitting, and yield more robust models [1, 29, 31, 32, 33, 34]. While there is an inherent trade-off between the number of ensemble members and the energy efficiency of a system [29, 31], the benefits of ensembling, particularly its enhanced accuracy and reliability, are invaluable to navigation-related robotics applications. Ensembling has been previously utilized in the context of monocular SLAM [35] for improved data association, terrain segmentation for learning data collected at different times [36], and place recognition with image-based [37, 38] and event-based [39] input, where multiple VPR techniques fuse to enhance place recognition robustness. Similarly, our work also uses ensembling to improve upon the robustness and generalization ability of our system but differs with these works in terms of how we define our ensemble members. Our ensemble members all are tasked to do the same task, have the same architecture and differ in random shuffling of input images and initial learned weights. A similar approach of employing ensembles was previously demonstrated in uncertainty estimation to provide sufficient diversity among the members, and improve the overall performance of the system [40].
In the context of VPR, a sequence matching technique uses consecutive reference frames instead of single frames, to match a query image of a place to its corresponding reference image [41, 42, 43]. By analyzing a series of images, this technique enables enhanced resilience against temporary environmental disruptions and improves localization accuracy [41, 42, 43]. Sequence matching techniques have been widely employed in place recognition [41, 44, 43, 45, 42, 46, 47], where often a decoupled approach consisting of an initial image-based retrieval method and subsequent sequence score aggregation [41, 44, 47] is employed. Similar to these techniques, our paper uses sequence matching as an additional step after image retrieval. We further demonstrate the effectiveness of sequence matching for SNN-based approaches, and provide an indicator for responsiveness of conventional and SNN-based VPR methods to sequence matching.
@@ II Related works
In this section, we offer an overview of neuromorphic computing and Spiking Neural Networks (SNNs) (Section II-A) and explore the applications of spiking neural networks in robot localization (Section II-B). We then delve into the Visual Place Recognition (VPR) task (Section II-C), non-spiking biologically inspired VPR approaches (Section II-D) and provide insights into ensembling (Section II-E) and sequence matching techniques (Section II-F).
@@ II-B Spiking Neural Networks for Robot Localization
The capabilities of SNNs have been demonstrated in a wide range of computer vision and robotics tasks. These include pattern recognition [73], control [74, 75, 76, 77, 78], manipulation [79, 80, 81], object tracking [82, 83], and scene understanding [84, 85]. Many works have employed SNNs to address tasks related to robot localization, which is the problem under investigation in this work. These works include computational models of place, grid and border cells of the rat hippocampus[86], a navigation controller for mapping unknown environments [87], a pose estimation and map formation method [88], a light-weight system for uni-dimensional SLAM [89], and a SLAM model that utilizes representations of continuous spatial maps to produce compressed structures from multiple domains [90]. In previous work [59], we presented a SNN specifically for VPR. This network had a limited capacity, recognizing up to just 100 places.
While some of these systems have been deployed on neuromorphic hardware [86, 88, 89, 91, 92], so far, the performance of SNN-based methods for robot localization have only been demonstrated in simulated environments [87, 88, 90], constrained indoor environments [86, 89, 91, 92, 93], or small-scale outdoor environments [59].
@@ II-C Visual Place Recognition
The VPR task is to determine whether a place has been previously visited, even when faced with appearance changes and perceptual aliasing [15, 17, 18, 14, 13, 20]. VPR is most-commonly framed as an image retrieval problem, where feature representations of a given query image are compared to the feature representations of all previously visited places, i.e. the reference database. The predicted place of the query image is the true position of the reference place that is most similar to the query in feature space. The VPR problem can also be posed as a template matching problem, similar to an image classification problem, where associated templates of all reference images are extracted to represent each place either via a single image [99, 49, 100, 48] or a sequence of images [41, 42, 101].
@@ II-E Ensembles of Neural Networks
One well-known approach to improve the predictive performance of neural networks is to use an ensemble of models [1, 29, 31, 32, 33, 34]. Ensembles have been shown to generalize well and prevent issues such as over-fitting and instability, which makes these approaches suitable to a wide range of applications in different domains [1, 29, 31, 32, 33, 34]. Challenges in deploying ensembles include requiring sufficient diversity in the output of the individual ensemble members, the trade-off between the computational complexity and the number of ensemble members, and the predictive performance and latency of the ensemble [1, 29, 31, 32, 33, 34]. Although ensembles are typically limited in scalability, our work anticipates leveraging neuromorphic computing, which has the potential to have highly efficient parallelization capability [8]. Consequently, neuromorphic deployment can alleviate scalability concerns in our ensembling approach due to the small size of each individual network.
Ensemble techniques have been used for a wide variety of robotic applications including image segmentation [114], gaze estimation [115], and uncertainty estimation [40]. In the field of SNNs, a variety of ensemble methods have been applied to diverse pattern recognition tasks [116, 117, 118, 119, 120, 121]. These methods include unsupervised ensembles of spiking expectation maximization networks [116], and ensembles of evolutionary SNN algorithms [119] that have been effective in digit recognition. Additionally, there are heterogeneous ensembles for few-shot online learning [117], and SNN ensembles that use a convolutional structure with unsupervised STDP learning [118], and ensembles of Liquid State Machines [121] which have been applied to image [117, 118, 121] and audio recognition [121] tasks. Furthermore, reservoir computing ensembles have been explored for multi-object behavior recognition [120].
@@ II-F Sequence Matching Techniques for Neural Networks
One common approach to improve the robustness of a VPR method, especially against sudden high appearance changes and perceptual aliasing, involves using the temporal information inherent in the database and query sets used in the mobile robot localization task [17, 15, 14, 13, 20]. One such set of algorithms that use the temporal information are sequence matching techniques, which can be classified into similarity-based methods, feature-based methods, and approaches that learn sequential information [13].
@@ III-A Preliminaries
Figure 2: SNN Module Architecture: Our Modular SNN is comprised by SNN modules that all have the same three-layer SNN architecture illustrated in this figure. Each module converts an input image to spike trains via rate coding, where the firing rate of input neurons is based on pixel intensities. The total number of input neurons is equal to the number of pixels in the input image, denoted as $K_{P}=W\times H$. These input neurons (blue dots) are fully connected to a layer of excitatory neurons (blue arrows connecting to green dots). Each excitatory neuron is connected to a single inhibitory neuron (single green arrow connecting to a red dot), which in turn connects to and inhibits all other excitatory neurons except its paired excitatory neuron (red arrows connecting back to green dots). The synaptic weights from excitatory to inhibitory neurons, $W_{EI}$, and from inhibitory back to excitatory neurons, $W_{IE}$, are fixed constants. The synaptic weights from input neurons to excitatory neurons, $W_{PE}$, are learned via the unsupervised Spike-Time Dependent Plasticity (STDP) mechanism that enables excitatory neurons to respond to different places. The number of output spikes from these excitatory neurons is used for place predictions.
@@ III-C Hyperactive Neuron Detection
The basic fusion approach, that considers all spiking neurons of all modules as presented in Eq. (7), is problematic. As the modules are only ever exposed to their local subset of the training data, there is a lack of global regularization to unseen training data outside of their local subset. In the case of spiking networks, this phenomenon leads to “hyperactive” neurons that are spuriously activated when stimulated with images from outside their training data. We decided to detect and remove these hyperactive neurons. This global regularization enhances place recognition capability within the modularity technique, allowing mapping to a larger number of places.
@@ III-D Ensemble of Modular SNNs
$D_{\text{single}}(u,v)=\max(S_{\text{single}})-S_{\text{single}}(u,v).$ (12)
@@ III-E Sequence Matching
where $I_{L_{seq}}$ is an identity matrix acting as a square filter kernel with dimensions of $L_{seq}\times L_{seq}$ pixels and $L_{seq}$ is the sequence length.
@@ IV-A Implementation Details
We implement our SNNs using the Brian2 simulator. We publicly release our code at: https://github.com/QVPR/VPRSNN. Each SNN module contains $K_{E}=400$ excitatory neurons, $K_{I}=400$ inhibitory neurons, and $K_{P}=W\times H$ input neurons, where the input image height and width is $28$ pixels in each dimension. We select the network’s hyperparameters as follows: For the number of training epochs, we use a fixed value of $30$ epochs for all datasets. For the threshold value $\theta$ (Eq. (8)) to detect the hyperactive neurons for each individual Modular SNN evaluated on a dataset, we draw the threshold value from a uniformly random distribution with $U(40,100)$ for small-scale datasets ($<1000$ places) and $U(600,800)$ for large-scale datasets ($>=1000$ places). The threshold value of each Modular SNN within an Ensemble of Modular SNNs evaluated on a dataset, is chosen separately based on this uniformly random selection process. For our Ensemble of Modular SNNs in the results section, Section V, we used three or five ensemble members, with the exact number specified in each part.
@@ IV-B Datasets
The Oxford RobotCar dataset [55]: has over 100 traverses captured in Oxford city and it is recorded at varying conditions including different time of the day, and different seasons. As done in prior works [131, 59], we selected the front left stereo frames from the Rain (2015-10-29-12-18-17) traverse as the reference dataset and from the Dusk (2014-11-21-16-07-03) traverse as the query dataset. For each traverse, we sampled approximately one image per meter, resulting in 3800 places.
The SFU-Mountain dataset [56]: has more than 8 hours of trail driving in Burnaby Mountain, British Columbia Canada, using the Clearpath Husky robot covering sunny, rainy and snowy conditions [56]. Following prior works [132, 133], we used the front right stereo frames from the dry traverse for reference and from the dusk traverse for query. We considered each image in the traverse as a different place, and used the author’s place sampling configurations [56], using the entire dataset where each traverse contains 375 places.
The Synthia Night-to-Fall dataset [57]: is a synthetic dataset that was initially designed for semantic scene understanding in a city-like driving scenario. In our approach, following the methodology described in [134], we used the SEQS-04 foggy (reference) and nighttime (query) traverses. Segments in which the vehicle remained stationary were excluded, and we sampled approximately one frame per meter resulting in 250 places for each traverse.
The St Lucia dataset [58]: comprises several traverses along a route within the St Lucia suburb of Brisbane. For our experiments, we employed a traverse conducted during the early morning (190809-0845) as the reference traverse and another in the afternoon (180809-1545) as the query traverse. We omitted the segments where the vehicle was at rest and sampled places approximately every 15 meters. We used only the unique places from the reference traverse, obtaining 500 places. For the query traverse, we included sections where places were visited multiple times within the same traverse, obtaining 1037 places.